OpenStudy (xapproachesinfinity):

find the maclaurin series for f(x) using known maclaurin series and evaluate \[f^{(4)}(0)\] \[f(x)=e^{\cos(2x)}\]

OpenStudy (xapproachesinfinity):

i tried to use maclaurin series for cos2x and take exp but seems bad tried to derived seems really tedious work with all those derivatives

OpenStudy (anonymous):

thank god for the internet, it is so smart http://www.sosmath.com/CBB/viewtopic.php?t=51552 scroll down to third answer

OpenStudy (xapproachesinfinity):

oh nice lol i didn't put on google to see.

OpenStudy (xapproachesinfinity):

that looks neat:)

OpenStudy (anonymous):

yours will be slightly different because of the \(2x\) but just as before the fact that it is even means there will only be even powers i have no idea what is means "use know maclaurin"

OpenStudy (xapproachesinfinity):

yeah i know sub will do :)

OpenStudy (amistre64):

ln(f(x)) = cos(2x) ln(f(x)) = sum (-1)^n (2x)^(2n)/(2n)! \[f(x)=exp\left(\sum_0 \frac{(-1)^n}{(2n)!} (2x)^{2n}\right)\] \[f'(x)= \left(\sum_1 \frac{(-1)^n4n}{(2n)!} (2x)^{2n-1}\right)~exp\left(\sum_0 \frac{(-1)^n}{(2n)!} (2x)^{2n}\right)\] \[f''(x)= \left(\sum_2 \frac{(-1)^n8n(2n-1)}{(2n)!} (2x)^{2n-2}+\sum_1 \frac{(-1)^n4n}{(2n)!} (2x)^{2n-1}\right)\\~~~~~~~~~~~~~~~~~~~~~~exp\left(\sum_0 \frac{(-1)^n}{(2n)!} (2x)^{2n}\right)\] pretty sure theres a pattern going here that can be simplified in a generality

OpenStudy (amistre64):

4 derivatives of the summations, and a multiple of the exp(...)

OpenStudy (xapproachesinfinity):

it crossed my mind to use logs

OpenStudy (amistre64):

with any luck \[f^{(4)}(x)= \\ \left( \sum_4 \frac{(-1)^n32n(2n-1)(2n-2)(2n-3)}{(2n)!} (2x)^{2n-4}\\ ~~~+\sum_3 \frac{(-1)^n16n(2n-1)(2n-2)}{(2n)!} (2x)^{2n-3}\\ ~~~+\sum_2 \frac{(-1)^n8n(2n-1)}{(2n)!} (2x)^{2n-2}\\ ~~~+\sum_1 \frac{(-1)^n4n}{(2n)!} (2x)^{2n-1}\right)\\ ~~~~~~~~~~~~~~~~~~~~~~exp\left(\sum_0 \frac{(-1)^n}{(2n)!} (2x)^{2n}\right)\]

OpenStudy (amistre64):

i dont think i see any constant in the summations x, x^2, x^3, x^4 ... at x=0 is 0 f^(4)(0) = 0

OpenStudy (xapproachesinfinity):

that's not the answer should be 64e

OpenStudy (amistre64):

yeah, 64e is the wolf .... do you see any error?

OpenStudy (amistre64):

cos(2x) power series is good

OpenStudy (xapproachesinfinity):

hmm seems good to me i think we need to align the indices there is a technique that we can use

OpenStudy (amistre64):

0:e^a 1:a' e^a 2:(a''+ a') e^a 3:(a'''+ a'' + a') e^a 4:(a''''+ a''' + a'' + a') e^a

OpenStudy (amistre64):

pulling out indexes just pulls out an x factor

OpenStudy (xapproachesinfinity):

i was doing a similar thing with u=a in my case could not keep track of the derivatives hahahA

OpenStudy (amistre64):

power series derivativs tend to be mechanical power rules ....

OpenStudy (xapproachesinfinity):

ok so a'=-2sin2x a''=-4cos2x a'''=8sin2x a''''=16cos2x f''''(x)=(16cos2x+8sin2x-4cos2x-2sin2x)e^cos2x f''''(0)=(16-4)e=12e

OpenStudy (xapproachesinfinity):

something is wrong with that hehe

OpenStudy (amistre64):

\[a=\sum_0 \frac{(-4)^n}{(2n)!} ~x^{2n}\implies 1-x^2+...\] \[a'=\sum_1 \frac{(-4)^n~2n}{(2n)!} ~x^{2n-1}\implies x+...\] \[a''=\sum_2 \frac{(-4)^n~2n(2n-1)}{(2n)!} ~x^{2n-2}\implies x^2+...\] \[a'''=\sum_3 \frac{(-4)^n~2n(2n-1)(2n-2)}{(2n)!} ~x^{2n-3}\implies x^3+...\] \[a''''=\sum_4 \frac{(-4)^n~2n(2n-1)(2n-2)(2n-3)}{(2n)!} ~x^{2n-4}\implies x^4+...\]

OpenStudy (xapproachesinfinity):

i think taking derivatives and ignoring taylor is good for this

OpenStudy (amistre64):

e^1 = e, but im getting 0s for the rest ... 1 + x^2 +x^4 + x^6 + x^8 2x +4x^3 + 6x^5 + 8x^7 2 +12x^2 + 30x^4 + 56x^6 24x +...

OpenStudy (amistre64):

i spose we should expect at 2 of the derivatives to have constants

OpenStudy (amistre64):

f'' and f''''

OpenStudy (xapproachesinfinity):

hmm I spent a lot of time on this this was given on the test can't believe it hahah

OpenStudy (xapproachesinfinity):

the prof is nuts, this needs at 2 hours by itself and there is another that is also challenging

OpenStudy (amistre64):

f^0 = 1 -2 x^2 + 2x^4/3 -4x^6/45 + ... f^1 = -4 x + 8x^3/3 -24x^5/45 + ... f^2 = -4 + 24x^2/3 -120x^4/45 + ... f^3 = 48x/3 -480x^3/45 + ... f^4 = 48/3 -480x^3/45 + ... (48/3 - 4) e (36/3) e 12e

OpenStudy (xapproachesinfinity):

should be 12e i keep getting the same thing too

OpenStudy (amistre64):

we have the power series right, we have the derivatives right by expansion .... we are centered at 0, which doesnt seem to give us a bad domain ....

OpenStudy (amistre64):

we know how to take the derivative of e^a, or e^u

OpenStudy (amistre64):

dont we? e^u u' e^u u'' e^u + u'u' e^u well, maybe we dont lol

OpenStudy (xapproachesinfinity):

i think we made a mistake with derivatives there lol

OpenStudy (amistre64):

yeah, looks that way

OpenStudy (xapproachesinfinity):

eh i give up i have others to finish

OpenStudy (xapproachesinfinity):

i will return to it and give another round

OpenStudy (amistre64):

we are almost done ... if we define the 4th derivative in terms of u ... we have the functon already derived and know some of its zero 2nd: u'' e^u + u'u' e^u 3rd: (u''' + u''u' + u'u'') e^u + (u'' + u'u')u' e^u 3rd: (u''' + 2u''u') e^u + (u'' u' + u'u'u') e^u 4th: (u'''' + 2u''' u'+ 2u'' u'') e^u + (u''' + 2u''u')u' e^u + (u''' u' + u'' u'' + 3u''(u')^2) e^u + (u'' u' + u'u'u')u' e^u u' = u''' = 0 (u'''' + 0+ 2u'' u'') e^u + (u''' + 2u''u')0 e^u + (0 + u'' u'' + 0) e^u + (u'' u' + u'u'u')0 e^u (u''''+ 3(u'')^2) e^u (48/3+ 3(4)^2) e (16+3(16)) e 4(16) e 64 e

OpenStudy (amistre64):

define u as a power series, so that you can define values of u,u',u'',u''',u'''' generalize the derivative process, and sub in afterwards ... thats the only way i would approach this

OpenStudy (xapproachesinfinity):

thanks a lot i will retry it myself:)